OpenStudy (anonymous):

the ratio of the lengths of the base and corresponding height of a triangle is 5 : #. if the area of the triangle is 270 millimeters squared, find the base and height.

OpenStudy (ajprincess):

What is the ratio @cassidykay?

OpenStudy (anonymous):

5 : 3

OpenStudy (ajprincess):

Let b be the base and h be the height. b:h=5:3 b/h=5/3 h=(3b)/5  Can u solve for b @cassidykay?